OpenStudy (anonymous):

X + Y = 48 875X + 1200Y =47,850 elimintaion method

OpenStudy (anonymous):

y=-x+48 875x + 1200y = 47,850 substitute the first equation into the second 875x + 1200(-x+48) = 47,860 875x -1200x + 57,600 = 47,860 -325X = -9740 x = 29.96